Rate Equation

Q1

a) Explain the meaning of the term rate of reaction (1 mark)

b) The following data is provided for the reaction

2NO(g) + H2(g) ——> N2O (g) + H2O

  • Initial (NO) (mol dm-3 ) 0.70 , Initial (H) ( mol dm-3) 0.42, Initial rate ( mol dm-3 s-1) 4.0 x 10-3.


  • Initial (NO) ( mol dm-3) 1.40, Initial (H) (mol dm-3) 0.42, initial rate ( mol dm-3 s-1) 1.6 x 10-2.


  • Initial (NO) (mol dm-3)1.40, Initial (H2) ( mol dm-3) 0.84, Initial rate (mol dm-3 s-1) 1.6 x 10-2.


(I)use the data to derive the order of reaction with respect to H and NO and explain your workings? (4 marks)

(II)Derive the overall order? (1 mark)

(III)Give the rate equation? (1 mark)

(IV)Calculate the value of k? (2 marks)


please detailed answers with expiation for everything please🙏🏾🙏🏾



1
Expert's answer
2022-01-17T17:26:02-0500

Part(a)


Rate of reaction or reaction rate is the speed at which a chemical reaction takes place, which indicates the amount of the reactant(s) consumed or the amount of the product(s) formed per unit time.


Part(b)

(i)

Let "Rate= k[NO]^m[H_2]^n"


Using data set 3 and 2


"\frac{1.6×10^{-2}}{1.6×10^{-2}}=\frac{k[1.4]^m[0.84]^n}{k[1.4]^m[0.42]^n}"


"1=2^n\,\,\implies n=0"


Using data set 2 and 1


"\frac{1.6×10^{-2}}{4×10^{-3}}=\frac{k[1.4]^m[0.42]^n}{k[0.7]^m[0.42]^n}"


"4=2^m\>\>\implies m=2"



The reaction is "0^{th}" order with respect to "H_2" and "2^{nd}" order with respect to "NO"


b(ii)

Overall order"= 0+2=2"


b(iii)

"Rate=k[NO]^2"


b(iv)

Using data set 1

"4\u00d710^{-3}=k[0.7]^2"


"k=8.163\u00d710^{-3} M^{-1}s^{-1}"